OpenStudy (anonymous):

what is the product in simplest form state any restrictions on the variable. 4a^5/7 b^4 * 2b^2/2a^4

OpenStudy (anonymous):

\[\frac{ 4a ^{5} }{ 7b ^{4} }*\frac{ 2b ^{2} }{ 2a ^{4} }\]

OpenStudy (anonymous):

See if you can cancel the higher exponents and then simplify.

OpenStudy (anonymous):

would i cancel out the 2s?

OpenStudy (anonymous):

yeah that is one thing you would do

OpenStudy (anonymous):

HINT: \[\frac{ a ^{5} }{ a ^{4} } = a\]

OpenStudy (anonymous):

can I cross simplify?

OpenStudy (anonymous):

\[2b^2/7b^4*2a?\]